# Stable soap bubble clusters with multiple torus bubbles

### v2

@@ 1 Abstract
In the last two centuries and more particularly in the last decades, the geometry of foams has become an important research domain, in mathematics, physics, material sciences and biology [4, 8, 16]. Most of the simplest geometrical observations of bubble clusters have long resisted rigorous mathematical proofs. Geometries can even get more complicated if immiscible fluids are considered. Although they have to fulfill Plateau’s laws [2, 10] like soap bubble clusters if the surface tensions are close to unity [9], this is not the case in general [8]. Whether we consider immiscible fluids clusters or soap type bubble clusters, an interesting question is the existence of stable clusters with topological torus bubbles. We propose to answer the latter numerically with simple numerical examples of stable clusters with bubbles having the topology of a torus, a triple torus, a fivefold torus or an elevenfold torus.
@@ 2 Introduction
Finding the geometry of a static foam consists in finding an energy minimizer (possibly local) of the bubble cluster for given volumes enclosed by each bubble. The energy is defined as the sum of the areas of the cluster interfaces. For immiscible fluids, each interface area can be weighted with a surface tension. One of the very first question about “foam” geometry is whether the smallest area enclosing a given volume is the sphere. This was formulated by Archimedes in the third century B.C.. This is only in the nineteenth century that it was proven to be true by Hermann A. Schwarz in 1884 [11]. The double bubble conjecture was formulated in the nineteenth century and states that the surface of minimal area enclosing two volumes is the standard double bubble, i.e. composed of three spherical caps sharing a circle intersection. The proof dates back from 2002 in [6]. The next natural question coming to mind is if larger clusters are also composed of spherical interfaces. In 2022, a proof [7] shows that the only bubble clusters of at most $3$ bubbles are the standard clusters, hence proving a conjecture of 1996 [14] (problem 2). It implies that clusters of at most $3$ bubbles only contain spherical interfaces. The problem for clusters of $4$ and $5$ bubbles remains open. For $6$ bubbles, an example [1, 8, 13] by John M. Sullivan exhibits a cluster containing a saddle shape interface hence shows that clusters of at least $6$ bubbles can contain non spherical interfaces. In fact, numerical experiments tend to show that large clusters are rarely composed of spherical interfaces exclusively and even rarely contain any spherical interface at all. Another interesting question of [14] (problem 1) that we propose to tackle here, is the existence of bubble clusters with bubbles being topologically a torus [^1] Note that torus bubbles have long been observed in the framework of dynamics [15].. We start with building a $5$-bubble cluster for immiscible fluids with a torus bubble. This cluster then serves as a basis to build stable soap bubble clusters with a triple, a fivefold or an elevenfold torus bubble. To do so, we duplicate the $5$-bubble cluster and assemble the copies around a tetrahedron, a cube or a dodecahedron.
@@ 5 Stability
The stability of the clusters has also been tested using the “longj” command (long jiggle) of Surface Evolver which moves the vertices according to a wavevector, a phase and a vector amplitude. After gradient steps and some mesh polishing, the clusters come back to their previous configurations. Though it is not its main purpose, perturbations with the “j” command have been tested as well. In this case, the vertices location is perturbed according to a Gaussian noise. The default “temperature” ($0.05$) as been chosen. Again, after some works, the clusters are back to their previous configurations.

### v3

@@ 1 Abstract
In the last two centuries and more particularly in the last decades, the geometry of foams has become an important research domain, in mathematics, physics, material sciences and biology [4, 8, 16]. Most of the simplest geometrical observations of bubble clusters have long resisted rigorous mathematical proofs. Geometries can even get more complicated if immiscible fluids are considered. Although they have to fulfill Plateau’s laws [2, 10] like soap bubble clusters if the surface tensions are close to unity [9], this is not the case in general [8]. In 1996, Frederick J. Almgren asked whether there is “any stable cluster of bubbles in $\mathbb{R}^{3}$ with some bubble being topolically a torus” [14] (problem 1). We propose to answer the latter numerically with simple numerical examples. We build stable soap bubble clusters with a triple torus bubble, a fivefold torus bubble or an elevenfold torus bubble. The construction uses the geometry of a simple immiscible fluids cluster with a torus bubble.
@@ 2 Introduction
Finding the geometry of a static foam consists in finding an energy minimizer (possibly local) of the bubble cluster for given volumes enclosed by each bubble. The energy is defined as the sum of the areas of the cluster interfaces. For immiscible fluids, each interface area can be weighted with a surface tension. One of the very first question about “foam” geometry is whether the smallest area enclosing a given volume is the sphere. This was formulated by Archimedes in the third century B.C.. This is only in the nineteenth century that it was proven to be true by Hermann A. Schwarz in 1884 [11]. The double bubble conjecture was formulated in the nineteenth century and states that the surface of minimal area enclosing two volumes is the standard double bubble, i.e. composed of three spherical caps sharing a circle intersection. The proof dates back from 2002 in [6]. The next natural question coming to mind is if larger clusters are also composed of spherical interfaces. In 2022, a proof [7] shows that the only bubble clusters of at most $3$ bubbles are the standard clusters, hence proving a conjecture of John M. Sullivan in 1996 [14] (problem 2). It implies that clusters of at most $3$ bubbles only contain spherical interfaces. The problem for clusters of $4$ and $5$ bubbles remains open. For $6$ bubbles, an example [1, 8, 13] by John M. Sullivan exhibits a cluster containing a saddle shape interface hence shows that clusters of at least $6$ bubbles can contain non spherical interfaces. In fact, numerical simulations tend to show that large clusters are rarely composed of spherical interfaces exclusively and maybe even rarely contain any spherical interface at all. Another interesting question by Frederick J. Almgren in 1996 [14] (problem 1) is whether there is “any stable cluster of bubbles in $\mathbb{R}^{3}$ with some bubble being topolically a torus” [^1] Note that torus bubbles have long been observed in the framework of dynamics [15].. We propose to tackle this problem in this document. We start with building a $5$-bubble cluster for immiscible fluids with a torus bubble. This cluster geometry is then used as a basis to build stable soap bubble clusters with a triple, a fivefold or an elevenfold torus bubble. To do so, we duplicate the $5$-bubble cluster and assemble the copies around a tetrahedron, a cube or a dodecahedron.
@@ 5 Stability
Although the command “longj” (long jiggle) is “archaic” according to Surface Evolver documentation, one can use it for a coarse study of the stability of the clusters. This command moves the vertices of the mesh according to a wavevector, a phase and a vector amplitude. Using random values provided by Surface Evolver, after gradient steps and some mesh polishing, the clusters are back to their previous stable configurations. Though it is not its main purpose, the “j” (jiggle) command can be tested as well. In this case, the vertices locations are perturbed according to a Gaussian noise. Using the default “temperature” ($0.05$), after some work as previously, the clusters are back to their previous stable configurations.
@@ 7 Conclusion
8 Acknowledgements
I would like to thank Professor Frank Morgan for pointing out mistakes in the abstract and introduction.